• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings

    Log entry time 08:09:59 on March13,2012

    Entry number 366198


    keyword=Owl Shift Summary

    Owl Shift Summary

    Shift Crew:
    SL: Melissa Cummings
    TO: Pete Markowitz
    3rd: Pengjia Zhu

    12:30 MCC sends tune beam into the hall.
    12:35 Pengjia starts his harp scans/BCM calibration
    01:00 Target NMR computer unresponsive - expert has been called in to fix the problem
    02:55 Hall C changed IHWP to IN
    03:15 Pengjia done with harp scans/raster calibration. Called for escorted access for sieve slit change. Target experts will try to fix target motion and network issues.
    03:45 Back in beam permit, beginning target polarization
    04:05 Alexandre is going to do some BCM tests
    04:45 Finished with BCM runs, target is at 68% polarization.
    04:55 Target is polarized to 69%, we still start taking some inelastic data
    05:30 Beam is not centered, moved target to empty holes so Pengjia could re-center the beam
    06:16 lost polarization, when beam comes back we will take data on the carbon target at this momentum setting
    06:20 Beam is still down - MCC called to say there is a problem at the injector and will probably be down for awhile.
    07:00 Called for restricted access. Ramping down target magnet.


    3152/22241 Po=2.050, NH3 target (bottom cup), IHWP is IN